Given the following Assumptions: You wish to earn $55000.00 90.00% of the MLS listings sell The average listing agents share is 1.5% of the selling price You plan to work 300 days next year The average sales price in the market is $90000.00 1 out of 4 phone calls will result in an appointment 1 out of 4 appointments will result in a listing How many telephone calls each day will you have to make to earn your desired income? (Express your answer as to two decimal places--X.xx) Your Answer:
Answer #1

Answer:

Earnings required = $55,000

Average listing agent's share is 1.5% of the selling price = Average sale price * agent's share = $90,000 * 1.5% = $1350

Number of listing required per annum = $55,000 /$1350 =40.74074

From all calls made, probability of getting a listing = Probability of getting an appointment out of all phone calls * Probability of getting a listing out of all appointments * Probability of MLS listing that sells

= 1/4 * 1/4 * 90%

= 5.625%

Hence total phone calls required to get 40.74074 listing = 40.74074 / 5.625% = 724.2798

Calls required per day = 724.2798 / 300 =2.41

2.41 calls per day